The number of telephone calls that arrive at a phone exchange is often modeled as a


Poisson random variable. Assume that on the average there are 10 calls per hour.


a) What is the probability that there are exactly 5 calls in one hour?


b) What is the probability that there are there are exactly 15 calls in two hours?


c) What is the probability that there are exactly 5 calls in 30 minutes?

1
Expert's answer
2022-03-25T14:29:37-0400

for Poisson distribution:

"P(x=k)= \frac{\lambda^ke^{-\lambda}}{k!}"

a. "P(5)=\frac{10^5e^{-10}}{5!}=0.0378"

b.The probability for 15 calls in 2 hours is "P(15,2)=\frac{(2\times 10)^{15}e^{-10\times 2}}{15!}=0.051"

c.The probability for 5 calls in 30 minutes or 0.5 hours is

"P(5,0.5)=\frac{(0.5\times10)^5e^{-10\times 0.5}}{5!}=0.175"